Find value of c so that the point (4,2) lies on the graph of the function f(x)=cx^2+3x-4
Answer by stanbon(75887)

Find value of c so that the point (4,2) lies on the graph of the function f(x)=cx^2+3x-4
-------------------
Substitute those values and solve for "c":
2 = c(4)^2 + 3*4 - 4
16c + 12 - 6 = 0
16c = -6
c = -3/8
================
Cheers,
Stan H.